# A text-to-tabular approach to generate synthetic patient data using LLMs

### v1

@@ 3.4 Evaluation framework of synthetic patient data generation
Fidelity metrics
To assess fidelity in terms of distribution shape, we calculated TVComplement for categorical variables (1 minus the Total Variation Distance (TVD), TVD is the distance between the frequency of each category), KSComplement for continuous variables (1 minus the Kolmogorov-Smirnov statistic, corresponding to the maximum difference between two estimated cumulative distribution functions (CDF)), and the weighted average of TVComplement and KSComplement, the Column Shape Score. Additionally, we computed two well-known distance metrics in the statistical field: the Wasserstein distance (WD) [Ramdas et al., 2015] and then Jensen-Shannon distance (JSD) [Lin, 1991].
@@
Privacy metrics
We used two distance metrics: distance to closest record (DCR) and nearest neighbour distance ratio (NNDR) [Lowe, 2004]. Following Zhao et al. [2021], to provide a robust estimate of privacy preservability, we computed the fifth percentile for both metrics. Additionally was also computed the percentage of new rows generated.
Moreover, we calculated a metric named CAPCategorical measuring privacy preservation against inference attacks using the gender, the MDS-UPDRS III and MoCA score as key fields known by the "attacker" and the number of years of education as a sensitive field.
Utility metric
@@
More detailed definitions of the computed metrics are provided in Appendix A. Additionally, since metrics can be challenging to interpret, our evaluation included univariate and joint bivariate distribution plots from the real and synthetic datasets.
3.5 Ablation study
We conducted several ablation studies on our prompt to better understand the key components contributing to the LLM’s performance. We conducted a comparison of several experiments on the PPMI data using fidelity metrics against a autoreference experiment. Specifically, we examined the impact on the quality of the synthetic data of using a newer and more advanced model (by switching from GPT-3.5 to GPT-4), of adding a single example patient to the prompt, and of sampling different number of synthetic patients $n$ at each iteration in the SDG process.
4 Results
4.1 Synthetic patient data that closely mirror the statistical properties of real data
The computed fidelity metrics showed that, overall, the synthetic cohorts preserved the real PD and AD cohorts’ statistical properties and patterns (respectively Table 1 and Table 2).
The synthetic data closely resembled the real PPMI and ADNI cohorts in terms of clinical characteristics. Indeed, the comparison of distribution plots between the synthetic and real data showed close resemblances in terms of distribution shape and parameters (Figure 3 and Figure 4 in Appendix B). However, the synthetic distributions had lower standard deviations and fewer outliers than the real data. This is likely due to the LLM generating data that closely mirror the most representative patients of PD and AD and to its temperature parameter.
SDG Model $D_{train}$ (Ref.) GPT-4 (Ours) GC CTGAN TVAE
@@ 4.1 Synthetic patient data that closely mirror the statistical properties of real data
Table 1: Benchmark results of SDG models on #PPMI2024 in terms of evaluation of fidelity, privacy preservation and utility (mean and standard deviation). For each metric, best results are reported in bold and second best results are underlined. Additionally is indicated the order of magnitude of the time to generate 1000 synthetic patients (SDG time). Wassersein distance (WD), Jenssen-Shannon distance (JSD), distance to closest record (DCR), nearest neighbour distance ratio (NNDR), Train Synthetic Test Real (TSTR), Train Augmented Test Real (TATR).
For both PPMI and ADNI cohorts, there was a noticeable shift in the mean and different distribution shapes of imaging variables: the left and right putamen variables for PPMI and the ventricles and intracranial volumes for ADNI (Figure 3 and Figure 4 in Appendix B). This gap could be explained by the sensitivity of these imaging variables to the PD and AD populations and by the poor available reported statistical analysis in the literature.
Moreover, the gender prevalence differences were notable and for PPMI the discrepancies for symptoms such as rigidity and tremor were even more pronounced when compared to the original data (Figure 3 in Appendix B). These inconsistencies might be explained from the broad range of prevalence values reported in the literature, as highlighted in the meta-analysis on the overall male/female prevalence ratios in PD by Zirra et al. [2022]. Furthermore, the subjective nature of the evaluation of symptoms, particularly rigidity, could contribute to these variations [Ferreira-Sánchez et al., 2020].
Finally, visual correlations between variables were well preserved (Figure 5 and Figure 6 in Appendix B).
4.2 An approach competing with established tabular-to-tabular approaches
Remarkably, despite not having access to the original data, our text-to-tabular approach achieved fidelity scores comparable to those of tabular-to-tabular approaches (Table 1). However, GC outperformed our method in terms of distribution shape and correlation.
SDG Model $D_{train}$ (Ref.) GPT-4 (Ours) Copula CTGAN TVAE
@@ 4.2 An approach competing with established tabular-to-tabular approaches
Table 2: Benchmark results of SDG models on #ADNI in terms of evaluation of fidelity, privacy preservation and utility (mean and standard deviation). For each metric, best results are reported in bold and second best results are underlined. Additionally is indicated the order of magnitude of the time to generate 1000 synthetic patients (SDG time). Wassersein distance (WD), Jenssen-Shannon distance (JSD), distance to closest record (DCR), nearest neighbour distance ratio (NNDR), Train Synthetic Test Real (TSTR), Train Augmented Test Real (TATR).
@@
Notably, no model exhibited significant overfitting on $D_{Train}$, as there were no substantial differences between scores computed on $D_{Train}$ and $D_{Test}$ (Table 8 and Table 2 in Appendix B). Our approach also excelled in preserving privacy, with the highest fifth percentile for DCR for PPMI, and the second-highest fifth percentile for NNDR after TVAE for both PPMI and ADNI (Table 1 and Table 2).
@@
4.3 Ablation study
Our ablation study showed that upgrading GPT-3.5 to GPT-4 improved the fidelity metrics (Table 3). Contrary to our expectation of an increase in data heterogeneity, sampling synthetic patients individually did not improve fidelity metrics.
@@ 4.3 Ablation study
autoreference GPT-3.5 10 0.64 0.84 0.58 0.88 0.94 0.79
Updated LLM GPT-4 10 0.71 0.83 0.68 0.80 0.95 0.75
Sampling 1 by 1 GPT-3.5 1 0.55 0.77 0.48 0.83 0.94 0.70
Example row added GPT-3.5 10 0.70 0.82 0.64 0.94 0.91 0.90
Best parameters GPT-4 10 0.76 0.82 0.74 0.81 0.94 0.75
Table 3: Ablation study
5 Discussion
This study highlights the potential of LLMs in generating synthetic healthcare data without access to the original data or requiring model pre-training or fine-tuning. Our text-to-tabular approach, relying solely on a description of the desired healthcare database and a frozen LLM, showed promising results by producing realistic synthetic tabular patient data. Applied to PPMI and ADNI datasets, our method successfully generated PD and AD patients’ data that preserved the characteristics of the distribution of the original data in spite of not having accessed it. Although our approach performed slightly worse in fidelity metrics compared to baseline SDG methods like GC, TVAE, and CTGAN, trained on the original data, it showed remarkable capabilities in replicating correlations between clinical variables, slightly outperforming CTGAN on some metrics. In addition, our method provides strong privacy guarantees by not using the original patient-level data.
Despite our efforts to provide a thorough evaluation of our approach, it is worth noting that there is currently no established consensus among researchers, regulatory bodies, and the pharmaceutical industry regarding the evaluation of synthetic patient data [Goncalves et al., 2020]. Evaluation should be application-specific and balance the trade-off between fidelity, utility and privacy [Alaa et al., 2022]. The level of evidence required to consider synthetic data reliable and validated for practical use should depend on the judgment criteria and the objective of use. Indeed, the evaluation criteria for synthetic data may vary depending on whether the data is used e.g., for data augmentation or for replacing control arms in regulatory submissions. Moreover, as highlighted in van Breugel et al. [2023], users should be careful and not treat synthetic data as real, particularly when applying it to ML tasks involving underrepresented groups, as generative models are likely to struggle with accurately modeling them.
While our approach demonstrates promising results, it has certain limitations. Firstly, our evaluation was conducted on two healthcare datasets focused on well-documented diseases and reference databases. It would be valuable to assess the performance of this approach on datasets with more restrictive access, other diseases, rare diseases, and of other nature such as longitudinal data. Extending the approach to other domains beyond healthcare and evaluating it on standard benchmark tabular datasets would also strengthen the validation.
@@ 5 Discussion
Appendix A Appendix
A.1 Data
A.1.1 PPMI
@@
Table 4: Attrition chart of PPMI study populations
A.1.2 ADNI
*Data used in preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). As such, the investigators within the ADNI contributed to the design and implementation of ADNI and/or provided data but did not participate in analysis or writing of this report. A complete listing of ADNI investigators can be found at: https://adni.loni.usc.edu/wp-content/uploads/how_to_apply/ADNI_Acknowledgement_List.pdf
@@ A.1.2 ADNI
Table 5: Attrition chart of ADNI study population
A.2 Data generation
@@ A.3.1 Fidelity
Univariate distributions plots
We compared distributions one by one of synthetic data against real data to examine the distribution shape and identify potential issues in the patterns learned by the SDG model.
Joint bivariate distribution plots
We compared scatter plots of pairs of continuous distributions to examine the relations that were maintained between the original and synthetic data.
@@
Column Shape Score: summary metric provided by the SDV library, is computed as the weighted average of the KSComplement and TVComplement scores Dat [2024].
Jensen-Shannon distance (JSD) Lin [1991]: distance computed as the square root of the Jensen-Shannon divergence, corresponding to a symmetrized and smoothed version of the Kullback–Leibler divergence; a measure of difference between two probability distributions. The distance ranges from 0 to 1 with 0 indicating that the two distributions are identical.
Wasserstein distance (WD) Ramdas et al. [2015]: also known as the earthmover distance, is a similarity metric that quantifies the amount of “work” (i.e. how much mass must be moved around, and how far) needed to transform one distribution into another. In 2D, it corresponds to the area between cumulative distribution functions. Values ranges between 0 and infinity, 0 corresponding to perfect overlap between the two distributions.
@@
A.3.2 Privacy preservation
@@
Distance to closest record (DCR): Distance to closest record measures the closest distance between a point in the synthetic data and the real data. The Euclidean distance is computed for each point of the synthetic data and then averaged. As suggested by Zhao et al. [2021], the 5th percentile of the metric was computed to provide a robust estimate of the privacy risk.
Nearest neighbor distance ratio (NNDR) Lowe [2004]: NNDR is a metric used to assess the proximity of a data point to outliers in the real data. It is defined as the ratio of the Euclidean distance to the nearest neighbor to the distance to the second nearest neighbor. An NNDR value close to 1 indicates that the point is likely situated within a cluster of the real data, whereas a value close to 0 suggests proximity to an outlier. By calculating NNDR for each synthetic record within the real data, we can identify whether the synthetic points are located in sparsely populated regions or within densely populated regions of the real data space. As well as for the DCR, the 5th percentile was provided for this metric.
A.3.3 Utility
@@
Appendix B Appendix
Name Description Group Category Type Mapping
moca MOCA Score (adjusted for education) Clinical assessment continuous int
DATSCAN_PUTAMEN_L Striatal Binding Ratio of the Left Putamen Small Volume Imaging continuous float
DATSCAN_PUTAMEN_R Striatal Binding Ratio of the Right Putamen Small Volume Imaging continuous float
updrs_totscore MDS-UPDRS Total Score (OFF) Clinical assessment continuous int
updrs1_score MDS-UPDRS Part I: Non-Motor Aspects of Experiences of Daily Living (OFF) Clinical assessment continuous int
updrs2_score MDS-UPDRS Part II: Motor Aspects of Experiences of Daily Living (OFF) Clinical assessment continuous int
updrs3_score MDS-UPDRS Part III: Motor Examination (OFF) Clinical assessment continuous int
duration Duration of Parkinson disease from Diagnosis to Visit Date Clinical assessment continuous float
EDUCYRS Years of Education Demographic continuous int
SEX Gender Demographic binary int {1: male, 0: female}
ageonset Age at onset of Parkinson disease diagnosis (Years) Demographic continuous float
sym_tremor Initial symptom (at diagnosis) - Resting Tremor Baseline symptom binary int {0: absence, 1: presence}
scopa SCOPA-AUT Total Score Clinical assessment continuous int
sym_rigid Initial symptom (at diagnosis) - Rigidity Baseline symptom binary int {0: absence, 1: presence}
Table 6: Description of a subset of variables from PPMI Clinical database (curated datacut v.20240129)
Name Description Group Category Type Mapping
AGE Age in years Baseline characteristic Categorical float
PTGENDER Sex Demographic Binary int {0: male, 1: female}
PTEDUCAT Years of education Demographic Continuous int
APOE4 APOE e4 carrier status Clinical assessment Continuous int
CDRSB CDR-SB Clinical assessment Continuous float
ADAS11 Alzheimer’s Disease Assessment Scale, 11-item version Clinical assessment Continuous float
MMSE Mini-Mental State Examination Clinical assessment Continuous float
Ventricles_bl UCSF Ventricles, volume of the brain’s ventricles measured by MRI, with the analysis conducted at the University of California, San Francisco (UCSF) Imaging Continuous int
WholeBrain_bl UCSF WholeBrain, volume of the whole brain measured by MRI, with the analysis conducted at the University of California, San Francisco (UCSF) Imaging Continuous int
ICV_bl UCSF ICV, intracranial volume measured by MRI, with the analysis conducted at the University of California, San Francisco (UCSF) Imaging Continuous int
Table 7: Description of a subset of variables from the ADNI dataset

### v2

@@ III-D1 Evaluation framework
III-D2 Fidelity metrics
To assess fidelity in terms of distribution shape, we calculated TVComplement for categorical variables (1 minus the Total Variation Distance (TVD), TVD is the distance between the frequency of each category), KSComplement for continuous variables (1 minus the Kolmogorov-Smirnov statistic, corresponding to the maximum difference between two estimated cumulative distribution functions (CDF)), and the weighted average of TVComplement and KSComplement, the Column Shape Score. Additionally, we computed two well-known distance metrics in the statistical field: the Wasserstein distance (WD) [39] and then Jensen-Shannon distance (JSD) [40].
@@ III-D2 Fidelity metrics
III-D3 Privacy metrics
We used two distance metrics: distance to closest record (DCR) and nearest neighbour distance ratio (NNDR) [41]. Following [19], to provide a robust estimate of privacy preservability, we computed the fifth percentile for both metrics. Additionally was also computed the percentage of new rows generated.
Moreover, we calculated a metric named CAPCategorical measuring privacy preservation against inference attacks using the gender, the MDS-UPDRS III and MoCA score as key fields known by the ”attacker” and the number of years of education as a sensitive field.
III-D4 Utility metric
@@
More detailed definitions of the computed metrics are provided in Appendix A. Additionally, since metrics can be challenging to interpret, our evaluation included univariate and joint bivariate distribution plots from the real and synthetic datasets (see Supplementary Materials).
III-E Ablation study
We conducted several ablation studies on our prompt to better understand the key components contributing to the LLM’s performance. We conducted a comparison of several experiments on the PPMI data using fidelity metrics against a reference experiment. Specifically, we examined the impact on the quality of the synthetic data of using a newer and more advanced model (by switching from GPT-3.5 to GPT-4), of adding a single example patient to the prompt, and of sampling different number of synthetic patients $n$ at each iteration in the SDG process.
IV Results
IV-A Synthetic patient data that closely mirror the statistical properties of real data
The computed fidelity metrics showed that, overall, the synthetic cohorts preserved the real PD and AD cohorts’ statistical properties and patterns (respectively Table I and Table II).
The synthetic data closely resembled the real PPMI and ADNI cohorts in terms of clinical characteristics. Indeed, the comparison of distribution plots between the synthetic and real data showed close resemblances in terms of distribution shape and parameters (Figure 5 and Figure 6 in Supplementary Materials). However, the synthetic distributions exhibited lower standard deviations and fewer outliers compared to the real data. These patterns likely reflect, respectively, the LLM’s tendency to mirror prototypical PD and AD patients, and the regularizing effect of the temperature parameter.
TABLE I: Benchmark results of SDG Models on #PPMI2024 in terms of evaluation of fidelity, privacy preservation and utility (mean and standard deviation). For each metric, best results are reported in bold and second best results are underlined. Additionally is indicated the order of magnitude of the time to generate 1000 synthetic patients (SDG time). Wassersein distance (WD), Jenssen-Shannon distance (JSD), distance to closest record (DCR), nearest neighbour distance ratio (NNDR), Train Synthetic Test Real (TSTR), Train Augmented Test Real (TATR).
SDG Model on #PPMI2024 $D_{train}$ (Ref.) GPT-4 (Ours) GC CTGAN TVAE
@@ IV-A Synthetic patient data that closely mirror the statistical properties of real data
For both PPMI and ADNI cohorts, there was a noticeable shift in the mean and different distribution shapes of imaging variables: the left and right putamen variables for PPMI and the ventricles and intracranial volumes for ADNI (Figure 5 and Figure 6 in Supplementary Materials). This gap could be explained by the sensitivity of these imaging variables to the PD and AD populations and by the poor available reported statistical analysis in the literature.
Moreover, the gender prevalence differences were notable and for PPMI the discrepancies for symptoms such as rigidity and tremor were even more pronounced when compared to the original data (Figure 5 in Supplementary Materials). These inconsistencies might be explained from the broad range of prevalence values reported in the literature, as highlighted in the meta-analysis on the overall male/female prevalence ratios in PD by [42]. Furthermore, the subjective nature of the evaluation of symptoms, particularly rigidity, could contribute to these variations [43].
Finally, visual correlations between variables were well preserved (Figure 7 and Figure 8 in Supplementary Materials).
IV-B An approach competing with established tabular-to-tabular approaches
Remarkably, despite not having access to the original data, our text-to-tabular approach achieved fidelity scores comparable to those of tabular-to-tabular approaches (Table I). However, GC outperformed our method in terms of distribution shape and correlation.
TABLE II: Benchmark results of SDG Models on #ADNI in terms of evaluation of fidelity, privacy preservation and utility (mean and standard deviation). For each metric, best results are reported in bold and second best results are underlined. Additionally is indicated the order of magnitude of the time to generate 1000 synthetic patients (SDG time). Wassersein distance (WD), Jenssen-Shannon distance (JSD), distance to closest record (DCR), nearest neighbour distance ratio (NNDR), Train Synthetic Test Real (TSTR), Train Augmented Test Real (TATR).
SDG Model on #ADNI $D_{train}$ (Ref.) GPT-4 (Ours) Copula CTGAN TVAE
@@ IV-B An approach competing with established tabular-to-tabular approaches
Notably, no model exhibited significant overfitting on $D_{Train}$, as there were no substantial differences between scores computed on $D_{Train}$ and $D_{Test}$ (Table VI for PPMI and Table IX in Supplementary Materials for ADNI). Our approach also excelled in preserving privacy, with the highest fifth percentile for DCR for PPMI, and the second-highest fifth percentile for NNDR after TVAE for both PPMI and ADNI (Table I and Table II).
@@
IV-C Ablation study
Our ablation study showed that upgrading GPT-3.5 to GPT-4 improved the fidelity metrics (Table V in Appendix A). Contrary to our expectation of an increase in data heterogeneity, sampling synthetic patients individually did not improve fidelity metrics.
@@ IV-C Ablation study
V Discussion
This study highlights the potential of LLMs in generating synthetic healthcare data without access to the original data and without requiring model fine-tuning. Our text-to-tabular approach, relying solely on a description of the desired healthcare database and a frozen LLM, showed promising results by producing realistic synthetic tabular patient data. Applied to PPMI and ADNI datasets, our method successfully generated PD and AD patients’ data that preserved the characteristics of the distribution of the original data in spite of not having accessed it. Although our approach performed slightly worse in fidelity metrics compared to baseline SDG methods like GC, TVAE, and CTGAN, trained on the original data, it showed remarkable capabilities in replicating correlations between clinical variables, slightly outperforming CTGAN on some metrics. In addition, our method provides strong privacy guarantees by not using the original patient-level data.
Despite our efforts to provide a thorough evaluation of our approach, it is worth noting that there is currently no established consensus among researchers, regulatory bodies, and the pharmaceutical industry regarding the evaluation of synthetic patient data [44]. Evaluation should be application-specific and balance the trade-off between fidelity, utility and privacy [45]. The level of evidence required to consider synthetic data reliable and validated for practical use should depend on the judgment criteria and the objective of use. Indeed, the evaluation criteria for synthetic data may vary depending on whether the data is used e.g., for data augmentation or for replacing control arms in regulatory submissions. Moreover, as highlighted in [46], users should be careful and not treat synthetic data as real, particularly when applying it to ML tasks involving underrepresented groups, as generative models are likely to struggle with accurately modeling them.
While our approach demonstrates promising results, it has certain limitations. Firstly, our evaluation was conducted on two healthcare datasets focused on well-documented diseases and reference databases. It would be valuable to assess the performance of this approach on datasets with more restrictive access, other diseases, rare diseases, and of other nature such as longitudinal data. To increase robustness of the validation, it would be valuable to extend the approach to a private benchmark of datasets or to custom data. Extending the approach to other domains beyond healthcare and evaluating it on standard benchmark tabular datasets would also strengthen the validation.
@@ V Discussion
Acknowledgment
Data collection and sharing for this project was funded by the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (National Institutes of Health Grant U01 AG024904) and DOD ADNI (Department of Defense award number W81XWH-12-2-0012). ADNI is funded by the National Institute on Aging, the National Institute of Biomedical Imaging and Bioengineering, and through generous contributions from the following: AbbVie, Alzheimer’s Association; Alzheimer’s Drug Discovery Foundation; Araclon Biotech; BioClinica, Inc.; Biogen; Bristol-Myers Squibb Company; CereSpir, Inc.; Cogstate; Eisai Inc.; Elan Pharmaceuticals, Inc.; Eli Lilly and Company; EuroImmun; F. Hoffmann-La Roche Ltd and its affiliated company Genentech, Inc.; Fujirebio; GE Healthcare; IXICO Ltd.; Janssen Alzheimer Immunotherapy Research & Development, LLC.; Johnson & Johnson Pharmaceutical Research & Development LLC.; Lumosity; Lundbeck; Merck & Co., Inc.; Meso Scale Diagnostics, LLC.; NeuroRx Research; Neurotrack Technologies; Novartis Pharmaceuticals Corporation; Pfizer Inc.; Piramal Imaging; Servier; Takeda Pharmaceutical Company; and Transition Therapeutics. The Canadian Institutes of Health Research is providing funds to support ADNI clinical sites in Canada. Private sector contributions are facilitated by the Foundation for the National Institutes of Health (www.fnih.org). The grantee organization is the Northern California Institute for Research and Education, and the study is coordinated by the Alzheimer’s Therapeutic Research Institute at the University of Southern California. ADNI data are disseminated by the Laboratory for Neuro Imaging at the University of Southern California.
VI Supplementary Materials
Supplementary materials are available online at https://github.com/quinten-health-os/synth-data-gen-from-text.
Appendix A
A-A Data
A-A1 PPMI
@@
TABLE III: Attrition chart of PPMI study populations
@@
A-A2 ADNI
Data used in preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). As such, the investigators within the ADNI contributed to the design and implementation of ADNI and/or provided data but did not participate in analysis or writing of this report. A complete listing of ADNI investigators can be found at: https://adni.loni.usc.edu/wp-content/uploads/how_to_apply/ADNI_Acknowledgement_List.pdf
TABLE IV: Attrition chart of ADNI study population
@@ A-A2 ADNI
A-B Data generation
@@ A-C1 Fidelity
  • Univariate distributions plots: We compared distributions one by one of synthetic data against real data to examine the distribution shape and identify potential issues in the patterns learned by the SDG Model.
  • Joint bivariate distribution plots: We compared scatter plots of pairs of continuous distributions to examine the relations that were maintained between the original and synthetic data.
@@
  • Column Shape Score: summary metric provided by the SDV library, is computed as the weighted average of the KSComplement and TVComplement scores [38].
  • Jensen-Shannon distance (JSD) [40]: distance computed as the square root of the Jensen-Shannon divergence, corresponding to a symmetrized and smoothed version of the Kullback–Leibler divergence; a measure of difference between two probability distributions. The distance ranges from 0 to 1 with 0 indicating that the two distributions are identical.
  • Wasserstein distance (WD) [39]: also known as the earthmover distance, is a similarity metric that quantifies the amount of “work” (i.e. how much mass must be moved around, and how far) needed to transform one distribution into another. In 2D, it corresponds to the area between cumulative distribution functions. Values ranges between 0 and infinity, 0 corresponding to perfect overlap between the two distributions.
@@
A-C2 Privacy preservation
  • Distance to closest record (DCR): Distance to closest record measures the closest distance between a point in the synthetic data and the real data. The Euclidean distance is computed for each point of the synthetic data and then averaged. As suggested by [19], the 5th percentile of the metric was computed to provide a robust estimate of the privacy risk.
  • Nearest neighbor distance ratio (NNDR) [41]: NNDR is a metric used to assess the proximity of a data point to outliers in the real data. It is defined as the ratio of the Euclidean distance to the nearest neighbor to the distance to the second nearest neighbor. An NNDR value close to 1 indicates that the point is likely situated within a cluster of the real data, whereas a value close to 0 suggests proximity to an outlier. By calculating NNDR for each synthetic record within the real data, we can identify whether the synthetic points are located in sparsely populated regions or within densely populated regions of the real data space. As well as for the DCR, the 5th percentile was provided for this metric.
@@ A-C2 Privacy preservation
A-C3 Utility
@@
A-D Ablation results
TABLE V: Ablation study
Reference GPT-3.5 10 0.64 0.84 0.58 0.88 0.94 0.79
Updated LLM GPT-4 10 0.71 0.83 0.68 0.80 0.95 0.75
Sampling 1 by 1 GPT-3.5 1 0.55 0.77 0.48 0.83 0.94 0.70
Example row added GPT-3.5 10 0.70 0.82 0.64 0.94 0.91 0.90
Best parameters GPT-4 10 0.76 0.82 0.74 0.81 0.94 0.75
Appendix B
Figure 3: Prompt used to generate a synthetic Parkinson’s disease population from PPMI.
TABLE VI: Benchmark results of SDG Models on #PPMI2024 in terms of evaluation of fidelity, privacy preservation and utility (mean and standard deviation). For each metric, the best results are reported in bold.